OpenStudy (anonymous):

sqrt(2p+4)-sqrt(p+3)=1

OpenStudy (anonymous):

\[\sqrt{p+4}-\sqrt{p+3}=1\]

OpenStudy (anonymous):

Is it 2p+4 or p+4 ?

OpenStudy (anonymous):

yes 2p...my mistake

OpenStudy (anonymous):

sqrt(2p+4)=1+sqrt(p+3) Squaring on both sides, 2p+4 = 1+p+3+2sqrt(p+3) , 2p+4 = 4+p+2sqrt(p+3) , p = 2sqrt(p+3). Again squaring on both sides, p^2 = 4(p+3) , p^2-4p-12=0 Solving the quadratic , p = -2,6.

OpenStudy (anonymous):

wouldnt it be -6 and 2??

OpenStudy (anonymous):

p^2-4p-12 factors to (p+6)(p-2) therefore p=-6,2 when i tried to chec my solutions 2 didnt work, but -6 did. so the correct solution is -6, correct?

OpenStudy (anonymous):

The factors are (p-6)(p+2). It will work .Please check it again.

OpenStudy (anonymous):

But the solution will be 6 only.

OpenStudy (anonymous):

ok, so i am correct -2 doesnt work? thank you

OpenStudy (anonymous):

You're welcome.